# Breaking Agent Backbones: Evaluating the Security of Backbone LLMs in AI Agents

### v1

@@ Abstract
AI agents powered by large language models (LLMs) are being deployed at scale, yet we lack a systematic understanding of how the choice of backbone LLM affects agent security. The non-deterministic sequential nature of AI agents complicates security modeling, while the integration of traditional software with AI components entangles novel LLM vulnerabilities with conventional security risks. Existing frameworks only partially address these challenges as they either capture specific vulnerabilities only or require modeling of complete agents. To address these limitations, we introduce threat snapshots: a framework that isolates specific states in an agent’s execution flow where LLM vulnerabilities manifest, enabling the systematic identification and categorization of security risks that propagate from the LLM to the agent level. We apply this framework to construct the $\operatorname{b}^{3}$ benchmark, a security benchmark based on 194331 unique crowdsourced adversarial attacks. We then evaluate 31 popular LLMs with it, revealing, among other insights, that enhanced reasoning capabilities improve security, while model size does not correlate with security. We release our benchmark, dataset, and evaluation code to facilitate widespread adoption by LLM providers and practitioners, offering guidance for agent developers and incentivizing model developers to prioritize backbone security improvements.
@@ 1 Introduction
In this paper, we aim to systematically understand how the choice of the backbone LLM in an AI agent affects its security. Many existing works have addressed similar questions from various perspectives. For instance, Debenedetti et al. (2024); Zhan et al. (2024); Liu et al. (2024); Zhang et al. (2025); Andriushchenko et al. (2025); Evtimov et al. (2025) all introduce benchmarks, competitions, or frameworks for evaluating the security of different types of AI agents. The frameworks employed within these works are, however, limited for two reasons: (i) The considered threats do not cover the full range of LLM vulnerabilities, e.g., by only considering indirect injections, remote code execution or other more restricted attack vectors. (ii) The frameworks require mocking entire agents including the full execution flow. This makes it both harder to convey the security implications and to achieve coverage across all threat types. Additionally, many works (e.g., Mazeika et al., 2024; Andriushchenko et al., 2025) focus on safety rather than security. In this paper, we distinguish between security and broader safety as follows: security concerns the ability of an adversary to exploit an agent in the context in which it is deployed. This is different from broader safety concerns around, e.g., toxicity and reliability.
@@
  • We combine the framework and data into the open-source $\operatorname{b}^{3}$ benchmark for LLM security, and use it to reveal actionable insights into the strengths and weaknesses of 31 popular backbone LLMs. Among other insights, our results reveal that enhanced reasoning capabilities improve security, while model size does not correlate with security.
@@ 2.2.1 A Threat Snapshot
Threat snapshots can be applied to several security-related tasks: (i) Threat modeling of a specific AI agent, which can then guide red-teaming engagements. In such a case, one can start by an exhaustive list of attack vectors and objectives in the agent and then build threat snapshots for each compatible pair. Even multi-step attacks, e.g., Crescendo attacks (Russinovich et al., 2025) or multi-agent attacks (Lee and Tiwari, 2024), can be modeled by decomposing them into multiple sequential threat snapshots. This may be challenging for a long-range attack, but we believe that even an imperfect decomposition of such an attack into individual threat snapshots provides in-depth understanding of the security vulnerabilities of the system. (ii) Benchmarking backbone LLMs, by creating threat snapshots that cover a broad range of use-cases, one can compare how different LLMs protect against threats. (iii) Contextual defenses: the maliciousness of payloads heavily depends on the context in which they are delivered. In this work, we only consider the second task and create a set of representative threat snapshots that can be used to benchmark backbone LLMs.
@@ 2.2.2 Attack categorization
An important part of threat snapshot modeling is to capture the full range of threats. This section presents our attack categorization, constructed specifically for this work rather than adapted from existing taxonomies, and demonstrates its broad coverage of threats affecting AI agents. We again restrict our discussion to attacks delivered in text form, but other modalities can be treated similarly. We propose two complementary categorizations: a vector-objective categorization based on attack vector and objective that facilitates threat modeling of individual AI agents; and a task-type categorization based on the affected LLM function that enables comparing fine-grained security properties of backbone LLMs.
@@ 3.1 Selecting Threat Snapshots
Our primary objective is to create a curated list of threat snapshots with comprehensive coverage across all attack categories both in terms of vector-objective and task-type (see Section 2.2.2), across fundamental differences in LLM content generation including tool-calling and structured outputs and across different ways of structuring the model context. We argue that having strong attacks is the most crucial component in getting a realistic assessment of the security of a threat snapshot. As a result, the list needs to be sufficiently small, so it is possible to collect strong attacks for each threat snapshot and run the benchmark efficiently.
@@ 3.2 Crowdsourcing Attack Collection
We collected high-efficacy targeted attacks through a closed beta trial of the Gandalf Agent Breaker challenge (Pfister et al., 2025), where users attempted to exploit our 10 threat snapshots described in Section 3.1. Each user was randomly assigned to one of 7 backbone LLMs (mistral-large, gpt-4o, gpt-4.1, o4-mini, gemini-2.5-pro, claude-3.7-sonnet, claude-3.5-haiku) [^3] All details including developer and API providers are given in Appendix E. and maintained this assignment across all attempts and difficulty levels to ensure consistent evaluation conditions.
@@
We recruited 947 users across 3 deployment waves to address early-stage platform issues. Based on user feedback during the trial, we refined threat snapshots to ensure consistent performance across backbone LLMs (e.g., reliable tool calling functionality). This iterative approach yielded a robust dataset of highly targeted, human-generated attacks for the representative set of threat snapshots. The final dataset of attacks contains 194331 unique attacks from 2400 player sessions [^4] A single player was assigned a new session if they cleared their browser cache or played on a new device., of which 10935 attacks were successful (above the score of 75 during the challenge).
@@ 4 Experiments
We evaluate a large list of 31 popular LLMs [^5] Details on the vendors and API providers that we use are given in Appendix E. on the $\operatorname{b}^{3}$ benchmark. Some of the selected models have configurable reasoning capabilities, so we evaluated them both with the reasoning disabled and enabled (2048 reasoning tokens, or medium reasoning effort, see Table 4 in Appendix C for the number of actually used reasoning tokens). Using the evaluation procedure described in Section 3.3 with $N=5$, we evaluate each model using the $210$ high-quality attacks selected in Section 3.2. This means that we collect 7 attacks per level and per snapshot, and submit the combined 21 attacks to each threat snapshot and level and repeat it 5 times to account for non-determinism in LLMs. The total vulnerability score for a given model is computed as in (1) and 95% bootstrap confidence intervals are computed as in (2).
@@ 4.1 Robustness of Attack Selection, Aggregation and Threat Snapshots
Figure 2: (top left) Model ranking is roughly preserved across key slices of threat snapshots, with some models standing out on tasks involving content safety and tool use. We plot models that perform the best or the worst in at least one category. (bottom left) LLMs with reasoning enabled have lower total vulnerability scores (lower is better). (right) Ranking based on total vulnerability scores for all models – lower score is better.
@@ 4.2 Benchmark Overall Ranking
We first consider the ranking based on the total vulnerability score (i.e., $\mathcal{T}$ consisting of all threat snapshots). It is provided in Figure 2 (right). The most secure models in our evaluation were grok-4, grok-4-fast and claude-opus-4-1 all with reasoning enabled. A more detailed analysis of the results provides several interesting insights.
@@
One of the most striking observations is that adding reasoning generally improves the security. Figure 2 (bottom left) directly compares the scores of all models for which reasoning can be enabled and disabled. As can be seen there is a clear improvement in the vulnerability scores for most models once reasoning is enabled. Interestingly, this contradicts conclusions drawn in (Zou et al., 2025). Additionally, we observe that the models for security decreases with reasoning are the tiny versions (i.e, gemini-2.5-flash-lite and gpt-5-nano), suggesting that reasoning requires a certain model size in order to bring an improvement.
@@
For most LLM benchmarks size has been a crucial indicator of performance. Interestingly, we did not observe a similar scaling behavior in our analysis. When comparing models for which different sizes were available (i.e., gpt-oss, llama4, gpt-5, claude-4 and gemini-2.5 based models), larger versions without reasoning capabilities showed no significant performance advantage over their smaller counterparts, and occasionally performed worse. When reasoning was enabled, we observed modest improvements with increased model size, though these gains were most pronounced in the transition from tiny to small variants (see Figure 5 in Appendix C for a direct comparison).
@@
Figure 2 (right) shows that top-rated systems use closed weights, indicating they are noticeably more secure than open-weight models. An important caveat: closed-weight systems typically incorporate additional guardrails and safety layers beyond the base model, whereas we evaluate open-weight models directly. This comparison therefore reflects system-level security for closed weights versus model-level security for open weights. Nevertheless, the best open weights model in our analysis is gpt-oss-120b with a score $0.44$ which is better than some of OpenAI’s older models (gpt-4o and gpt-4.1) and only slightly worse than the current gpt-5 based models. This indicates that while there is a performance gap, open weights models are not lagging too far behind.
Effects of release date and price.
While newer models appear to perform slightly better, the improvement in security is actually not as substantial as one might expect (see Figure 7 in Appendix C). Secondly, more expensive models generally rank higher (see Figure 8 (right) in Appendix C).
@@ 4.3 Benchmark Ranking Based on Task Type and Other Categories
Our benchmark allows us to rank based on sub-categories as discussed in Section 3.3. The full results for the rankings based on categories is shown in Figure 2 (top left) and for defenses and task types in Figures 10 and 11 in Appendix C, respectively. We draw the following conclusions.
@@
grok-4 remains the most secure across each individual defense layer $\operatorname{L1}$, $\operatorname{L2}$ and $\operatorname{L3}$ (see Figure 10 in Appendix C). This suggests that the employed type of defense used should not influence LLM backbone selection. In particular, the worst and the best models perform consistently in $\operatorname{L1}$ and $\operatorname{L2}$, which differ only by prompt strength.
LLM security is different from safety.
Different models take the lead when tasks involve content safety abuse, see Figure 2 (top left). For example, gemini-2.5-pro with reasoning, which takes the 13th place in the full ranking, takes the lead in the sub-ranking for content safety.
@@
When looking at model performance split by defense levels or threat snapshot categories, the best and worst models’ performance is relatively consistent. That is not the case when slicing the results by task types, as seen in Figure 11 in Appendix C. This fact highlights that a model’s security properties differ between task types, and thus backbones should be chosen with a specific use case in mind.
@@ 5 Discussion
We defined and isolated the novel vulnerability affecting LLMs and introduced the threat snapshot framework as a corresponding abstraction. By presenting a corresponding exhaustive attack categorization, we created threat snapshots with broad coverage of LLM vulnerabilities within AI agents and collected a high-quality crowdsourced adversarial attack dataset. We then combined these ingredients to create the $\operatorname{b}^{3}$ benchmark and used it to evaluate 31 popular LLMs. Surprisingly, features such as an LLM’s size do not correlate with its security. Importantly, by considering subsets of threat snapshots among key dimensions, the $\operatorname{b}^{3}$ benchmark provides more fine-grained insights unavailable in existing security benchmarks. These findings provide actionable insights for developers to select the most secure backbone LLM for their specific agentic use case.
@@
There are, however, some limitations of our work. First, we did not consider utility or latency of any of the models. Other benchmarks exist for such purposes and selecting a backbone should take the resulting security-utility tradeoff into account. For example, a coding agentic use case should additionally consider the performance of the backbone LLM in coding benchmarks. Second, we focused only on evaluating the backbone LLMs in AI agents. Future work could extend this and apply the threat snapshot framework to red-teaming or evaluating external defenses deployed in agentic systems. Finally, backbone security is only a component of agentic security. While threat snapshots allow us to model the effect of, for example, a poisoned document on the backbone LLM, it does not allow us to quantify the likelihood that such a document is retrieved from a given RAG implementation in the first place. Further work should focus on how attacks propagate beyond a single step in the agentic flow and how they interact with other software components in the system.
Acknowledgments
We would like to thank all of the participants that took part beta trial of the Gandalf Agent Breaker challenge.
@@ Acknowledgments
  • Appendix B: Details on Selected Threat Snapshots

### v2

@@ Abstract
AI agents powered by large language models (LLMs) are being deployed at scale, yet we lack a systematic understanding of how the choice of backbone LLM affects agent security. The non-deterministic sequential nature of AI agents complicates security modeling, while the integration of traditional software with AI components entangles novel LLM vulnerabilities with conventional security risks. Existing frameworks only partially address these challenges as they either capture specific vulnerabilities only or require modeling of complete agents. To address these limitations, we introduce threat snapshots: a framework that isolates specific states in an agent’s execution flow where LLM vulnerabilities manifest, enabling the systematic identification and categorization of security risks that propagate from the LLM to the agent level. We apply this framework to construct the $\operatorname{b}^{3}$ benchmark, a security benchmark based on 194,331 unique crowdsourced adversarial attacks. We then evaluate 34 popular LLMs with it, revealing, among other insights, that enhanced reasoning capabilities improve security, while model size does not correlate with security. We release our benchmark, dataset, and evaluation code to facilitate widespread adoption by LLM providers and practitioners, offering guidance for agent developers and incentivizing model developers to prioritize backbone security improvements.
@@ 1 Introduction
In this paper, we aim to systematically understand how the choice of the backbone LLM in an AI agent affects its security. Many existing works have addressed similar questions from various perspectives. For instance, Debenedetti et al. (2024); Zhan et al. (2024); Liu et al. (2024); Zhang et al. (2025); Andriushchenko et al. (2025); Evtimov et al. (2025) all introduce benchmarks, competitions, or frameworks for evaluating the security of different types of AI agents. The frameworks employed within these works are, however, limited for two reasons: (i) The considered threats do not cover the full range of LLM vulnerabilities, e.g., by only considering indirect injections, remote code execution or other more restricted attack vectors. (ii) The frameworks require mocking entire agents including the full execution flow. A detailed comparison to existing agent security benchmarks is provided in Appendix G. This makes it both harder to convey the security implications and to achieve coverage across all threat types. Additionally, many works (e.g., Mazeika et al., 2024; Andriushchenko et al., 2025) focus on safety rather than security. In this paper, we distinguish between security and broader safety as follows: security concerns the ability of an adversary to exploit an agent in the context in which it is deployed. This is different from broader safety concerns around, e.g., toxicity and reliability.
@@
  • We combine the framework and data into the open-source $\operatorname{b}^{3}$ benchmark for LLM security, and use it to reveal actionable insights into the strengths and weaknesses of 34 popular backbone LLMs. Among other insights, our results reveal that enhanced reasoning capabilities improve security, while model size does not correlate with security.
@@ 2.2.1 A Threat Snapshot
Threat snapshots can be applied to several security-related tasks: (i) Threat modeling of a specific AI agent, which can then guide red-teaming engagements. In such a case, one can start by an exhaustive list of attack vectors and objectives in the agent and then build threat snapshots for each compatible pair. (ii) Benchmarking backbone LLMs, by creating threat snapshots that cover a broad range of use-cases, one can compare how different LLMs protect against threats. (iii) Contextual defenses: the maliciousness of payloads heavily depends on the context in which they are delivered.
While threat snapshots by design capture single LLM calls, the framework also applies to multi-turn (e.g., Crescendo attacks (Russinovich et al., 2025)) and multi-agent attacks (Lee and Tiwari, 2024). Such multi-step threats can be modeled by decomposing a concerning outcome into a chain of threat snapshots, where each threat snapshot represents one required step toward that outcome. This decomposition is valid because LLMs are stateless – each call receives full context needed for inference – making threat snapshots a complete atomic abstraction regardless of complexity. We provide two example decompositions in Appendix B.2.
In this work, we consider task (ii) and focus only on single-step threat snapshots in order to cover sufficiently diverse threats to benchmark backbone LLMs across a broad range of application scenarios.
@@ 2.2.2 Attack categorization
An important part of threat snapshot modeling is to capture the full range of threats. This section presents our attack categorization, constructed specifically for this work rather than adapted from existing taxonomies, and demonstrates its broad coverage of threats affecting AI agents. The agent perspective is necessary to systematically cover threats beyond existing LLM prompt-injection benchmarks, including tool-related attack vectors (e.g., poisoned tool definitions), tool invocation task types (discussed below), and attack objectives such as system and tool compromise. We again restrict our discussion to attacks delivered in text form, but other modalities can be treated similarly. We propose two complementary categorizations: a vector-objective categorization based on attack vector and objective that facilitates threat modeling of individual AI agents; and a task-type categorization based on the affected LLM function that enables comparing fine-grained security properties of backbone LLMs.
@@ 3.1 Selecting Threat Snapshots
Our primary objective is to create a curated list of threat snapshots with comprehensive coverage across all attack categories both in terms of vector-objective and task-type (see Section 2.2.2), across fundamental differences in LLM content generation including tool-calling and structured outputs and across different ways of structuring the model context. We argue that having strong attacks is the most crucial component in getting a realistic assessment of the security of a threat snapshot. As a result, the list needs to be sufficiently small, so it is possible to collect strong attacks for each threat snapshot and run the benchmark efficiently. We focus on single-step attacks since LLMs weak in single-step settings will be weak in multi-step scenarios, establishing a lower bound on system security. This simplified scope enables broader coverage across attack categories and more fine-grained assessment.
@@
While our attack categorization aims to be comprehensive, we acknowledge that, as with any security benchmark, blindspots may exist, whether from novel ways to exploit backbone LLM vulnerabilities or new attack surfaces emerging as agent architectures evolve. The modular threat snapshot framework, however, makes it straightforward to extend the benchmark as new threats are identified.
@@ 3.2 Crowdsourcing Attack Collection
We collected high-efficacy targeted attacks through a closed beta trial of the Gandalf Agent Breaker challenge (Pfister et al., 2025), where users attempted to exploit our 10 threat snapshots described in Section 3.1. Each user was randomly assigned to one of 7 backbone LLMs (mistral-large, gpt-4o, gpt-4.1, o4-mini, gemini-2.5-pro, claude-3-7-sonnet, claude-3-5-haiku) [^3] All details including developer and API providers are given in Appendix E. and maintained this assignment across all attempts and difficulty levels to ensure consistent evaluation conditions.
@@
We recruited 947 users across 4 deployment waves to address early-stage platform issues. Based on user feedback during the trial, we refined threat snapshots to ensure consistent performance across backbone LLMs (e.g., reliable tool calling functionality). This iterative approach yielded a robust dataset of highly targeted, human-generated attacks for the representative set of threat snapshots. The final dataset of attacks contains 194,331 unique attacks from 13,920 player sessions [^4] A single player was assigned a new session if they cleared their browser cache or played on a new device., of which 10,935 attacks were successful (above the score of 75 during the challenge).
@@ 4 Experiments
We evaluate a large list of 34 popular LLMs [^5] Details on the vendors and API providers that we use are given in Appendix E. on the $\operatorname{b}^{3}$ benchmark. Some of the selected models have configurable reasoning capabilities, so we evaluated them both with the reasoning disabled and enabled (2048 reasoning tokens, or medium reasoning effort, see Table 4 in Appendix C for the number of actually used reasoning tokens). Using the evaluation procedure described in Section 3.3 with $N=5$, we evaluate each model using the $210$ high-quality attacks selected in Section 3.2. This means that we collect 7 attacks per level and per snapshot, and submit the combined 21 attacks to each threat snapshot and level and repeat it 5 times to account for non-determinism in LLMs. The total vulnerability score for a given model is computed as in (1) and 95% bootstrap confidence intervals are computed as in (2).
@@ 4.1 Robustness of Attack Selection, Aggregation and Threat Snapshots
Figure 2: (top left) Vulnerability scores for each task type (see Section 2.2.2), showing that the security of a model depends on the task type. We only include models that perform the best or the worst in at least one task type. (bottom left) LLMs with reasoning enabled have lower total vulnerability scores (lower is better). (right) Ranking based on total vulnerability scores for all models – lower score is better.
@@ 4.2 Benchmark Overall Ranking
We first consider the ranking based on the total vulnerability score (i.e., $\mathcal{T}$ consisting of all threat snapshots). It is provided in Figure 2 (right). The most secure models in our evaluation were claude-haiku-4-5, claude-sonnet-4-5 with reasoning enabled and claude-haiku-4-5 without. A more detailed analysis of the results provides several interesting insights.
@@
One of the most striking observations is that adding reasoning generally improves the security. Figure 2 (bottom left) directly compares the scores of all models for which reasoning can be enabled and disabled. As can be seen there is a clear improvement in the vulnerability scores for most models once reasoning is enabled. Interestingly, this contradicts conclusions drawn in (Zou et al., 2025). Moreover, only the tiny model versions exhibit decreased security as reasoning increases (i.e, gemini-2.5-flash-lite and gpt-5-nano), suggesting that reasoning requires a certain model size in order to bring an improvement.
@@
For most LLM benchmarks size has been a crucial indicator of performance. Interestingly, we did not observe a similar scaling behavior in our analysis. When comparing models for which different sizes were available (i.e., gpt-oss, llama4, gpt-5, claude-4, claude-4-5 and gemini-2.5 based models), larger versions without reasoning capabilities showed no significant performance advantage over their smaller counterparts, and occasionally performed worse. When reasoning was enabled, we observed modest improvements with increased model size, though these gains were most pronounced in the transition from tiny to small variants (see Figure 9 in Appendix C for a direct comparison).
@@
Figure 2 (bottom right) shows that top-rated systems use closed weights, indicating they are noticeably more secure than open-weight models. An important caveat: closed-weight systems typically incorporate additional guardrails and safety layers beyond the base model, whereas we evaluate open-weight models directly. This comparison therefore reflects system-level security for closed weights versus model-level security for open weights. Nevertheless, the best open weights model in our analysis is kimi-k2-thinking with a score $0.34$ which is better than OpenAI’s frontier gpt-5.1 model (without reasoning). This indicates that while there is a performance gap, open weights models are not lagging too far behind.
Capability correlates with security but with exceptions
Newer more capable models exhibit superior security compared to their predecessors (Figure 11, Appendix C), likely due to improved instruction-following that reduces confusion between system and external instruction and able to better understand context. We confirm this hypothesis by comparing vulnerability scores against the agentic intelligence index (Artificial Analysis, 2025) (Figure 16, Appendix C), which shows a clear correlation between capability and security. Notably, however, the two highest-capability models (gpt-5.1 and kimi-k2-thinking) rank only 8th and 14th in security, indicating that capability alone is insufficient for security.
@@ 4.3 Benchmark Ranking Based on Task Type and Other Categories
Our benchmark allows us to rank based on sub-categories as discussed in Section 3.3. The full results for the rankings based on categories is shown in Figure 2 (top left) and for defenses and categories in Figures 14 and 15 in Appendix C, respectively. We draw the following conclusions.
@@
claude-haiku-4-5 remains the most secure across each individual defense layer $\operatorname{L1}$, $\operatorname{L2}$ and $\operatorname{L3}$ (see Figure 14 in Appendix C). This suggests that the employed type of defense used should not influence LLM backbone selection. In particular, the worst and the best models perform consistently in $\operatorname{L1}$ and $\operatorname{L2}$, which differ only by prompt strength.
@@
When looking at model performance split by defense levels or threat snapshot categories, the best and worst models’ performance is relatively consistent. That is not the case when slicing the results by task types, as seen in Figure 2 (top left) in Appendix C. This fact highlights that a model’s security properties differ between task types, and thus backbones should be chosen with a specific use case in mind.
@@ 5 Discussion
We defined and isolated the novel vulnerability affecting LLMs and introduced the threat snapshot framework as a corresponding abstraction. By presenting a corresponding exhaustive attack categorization, we created threat snapshots with broad coverage of LLM vulnerabilities within AI agents and collected a high-quality crowdsourced adversarial attack dataset. We then combined these ingredients to create the $\operatorname{b}^{3}$ benchmark and used it to evaluate 34 popular LLMs. Surprisingly, features such as an LLM’s size do not correlate with its security. Importantly, by considering subsets of threat snapshots among key dimensions, the $\operatorname{b}^{3}$ benchmark provides more fine-grained insights unavailable in existing security benchmarks. These findings provide actionable insights for developers to select the most secure backbone LLM for their specific agentic use case.
@@
There are, however, some limitations of our work. First, we did not consider utility (at their intended tasks) or latency of any of the models, as we believe the benchmarking focus to date has mostly focused on utility already. Other benchmarks exist for such purposes and selecting a backbone should take the resulting security-utility tradeoff into account as proposed in the D-SEC framework (Pfister et al., 2025). For example, a coding agentic use case should additionally consider the performance of the backbone LLM in coding benchmarks. We show how $\operatorname{b}^{3}$ scores compare to a general utility benchmark in Figure 14. Second, we focused only on evaluating the backbone LLMs in AI agents. Future work could extend this and apply the threat snapshot framework to red-teaming or evaluating external defenses deployed in agentic systems. Third, backbone security is only one component of agentic security. While threat snapshots allow us to model the effect of, for example, a poisoned document on the backbone LLM, it does not allow us to quantify the likelihood that such a document is retrieved from a given RAG implementation in the first place. Further work should focus on how attacks propagate beyond a single step in the agentic flow and how they interact with other software components in the system. Finally, crowdsourcing attacks is resource-intensive and could run into limitations once AI systems become sufficiently powerful. Automated attack generation methods could improve scalability and reduce bias. While current automated attacks are insufficient for our context-dependent threat snapshots, our framework is designed to support future integration with reinforcement learning and iterative methods, enabling the development of adaptive attacks that co-evolve with model capabilities.
Ethical Statement
Our research involves the collection and public release of a large-scale dataset of adversarial attacks against LLMs, crowdsourced from human volunteers. We acknowledge the dual-use nature of releasing attack data, but emphasize that the attacks are highly targeted to corresponding threat snapshot contexts. As such it is non-trivial to transfer these attacks to new settings. We show that some transfer is possible in Section 4.1, however, the quality of the attacks decreases substantially as part of this transfer. To further mitigate any misuse, we only plan to publish the lower quality version of the attacks for which the most effective attacks have been removed. We believe the security benefits of enabling widespread defensive improvements substantially outweigh the risks of potential misuse. Finally, prior to release, we are contacting all affected LLM providers and giving them the option of patching their models before releasing the data.
Reproducibility Statement
We integrated the benchmark within the Inspect framework (AI Security Institute, UK, 2024) and created a public code repository https://ukgovernmentbeis.github.io/inspect_evals/evals/safeguards/b3/ that contains all required code to run the benchmark. The repository includes all of the threat snapshots and the modified threat snapshots used in Section 4.1. A public version of the lower quality attack dataset is available at https://huggingface.co/datasets/Lakera/b3-agent-security-benchmark-weak.
Disclosure of LLM Use
We used LLMs exclusively for two purposes while writing the paper: (i) light editing, such as minor rephrasings and grammatical checks; all substantive content and analysis were done by the authors and (ii) help in generating and prettifying some of the plots in the paper.
@@
  • Appendix B: Details on Threat Snapshots
@@
  • Appendix G: Comparison to other Agent Security Benchmarks